# Building, Reusing, and Generalizing Abstract Representations from Concrete Sequences

### v1

@@ Abstract
Humans excel at learning abstract patterns across different sequences, filtering out irrelevant details, and transferring these generalized concepts to new sequences. In contrast, many sequence learning models lack the ability to abstract, which leads to memory inefficiency and poor transfer. We introduce a non-parametric hierarchical variable learning model (HVM) that learns chunks from sequences and abstracts contextually similar chunks as variables. HVM efficiently organizes memory while uncovering abstractions, leading to compact sequence representations. When learning on language datasets such as babyLM, HVM learns a more efficient dictionary than standard compression algorithms such as Lempel-Ziv. In a sequence recall task requiring the acquisition and transfer of variables embedded in sequences, we demonstrate HVM’s sequence likelihood correlates with human recall times. In contrast, large language models (LLMs) struggle to transfer abstract variables as effectively as humans. From HVM’s adjustable layer of abstraction, we demonstrate that the model realizes a precise trade-off between compression and generalization. Our work offers a cognitive model that captures the learning and transfer of abstract representations in human cognition and differentiates itself from the behavior of large language models.
@@ 1 Introduction
We hypothesize that the world contains patterns across scales of time and abstraction. Intelligent agents-facing sequences with nested hierarchical structures- need to model these structures to store, process, and interact in such environments. As a rational strategy, intelligent agents shall characterize the temporal structure via chunking and characterize the abstract structure via identifying common features and grouping collections of different items that play similar roles. To explore these operations, we design a generative model that produces sequences nested with hierarchical structure and propose an approximate recognition model that conjunctively learns to chunk and abstract.
@@
We go beyond previous proposals that introduce chunking as a mechanism for learning and composing complex structures from elementary perceptual units (Gobet et al., 2001; Miller, 1956; Wu et al., 2022) and propose a cognitive model that combines chunking and abstraction in one single system. The model uses abstraction in two key ways: first, by identifying shared features to facilitate efficient pattern retrieval, such as recognizing patterns from learned animals when fur is observed; and second, by categorizing different sequential items that appear in the same context, much like a variable in a computer program— for example, using the category "animal" to represent a cat, dog, or squirrel. These paired mechanisms enable the model to parse sequences into chunks and form abstract patterns based on both concrete and previously learned abstractions. Abstraction allows for memory-efficient, compact representations and reveals recurring patterns at the level of abstract categories. This allows the model to discover increasingly complex or abstract patterns, layer by layer, with controllably varying degrees of detail.
@@
We first demonstrate the benefits of abstraction in memory efficiency and sequence parsing by comparing our algorithm with previous chunking models and other dictionary-based compression methods. Then, we show that the model exhibits human-like signatures of abstraction in a memory experiment requiring the transfer of abstract concepts. In the same experiment, we contrast the model’s generalization behavior with large language models (LLMs), which have demonstrated rudimentary reasoning abilities (Wei et al., 2022) in recent years. We show that LLMs do not transfer abstract variables and rely more on associative learning than abstraction. We demonstrate the connection between abstraction level and abstract concept transfer by varying the level of abstraction as a parameter in the model. Our work offers a cognitive model that captures the learning and transfer of abstract representations in human cognition and differentiates itself from the behavior of artificial agents.
@@
Figure 1: Generative Model. The inventory is initialized with a set of atomic units. These units randomly combine into objects. Some objects are randomly selected to become categories. Categories of existing objects have similar interaction properties. The dashed arrow points from those objects to their corresponding categories. These categories further combine with existing objects. An observational sequence is composed of objects randomly sampled from the existing inventory. Categories are latent and never explicitly observed but are manifested in one of the denoting objects.
@@ 2 Generating sequences with objects that contain hierarchical abstract structures
We design a probabilistic generative model that generates sequences by sampling objects from an inventory, with these objects organized hierarchically, resembling the structure found in natural sequences—such as molecules composed of chemical elements or the hierarchical nature of language (Abler, 1989; Sportiche et al., 2013; von Humboldt & Losonsky, 1999).
The generative model creates an inventory of recurring objects over $d$ iterations of expansion. As illustrated in Figure 1, the inventory starts with a set of atomic units $\mathbb{A}$. On each iteration, a novel object or category is created equiprobably. A new category (graded node) is created by pointing to a random selection of objects from the inventory up to the moment (these objects are treated disjunctively). A new object is created by concatenating a random selection of pre-existing objects or categories from the existing inventory. After the inventory has been expanded up to the $d$-th iteration, the independent occurrence probability in the sequence is sampled from a flat Dirichlet distribution $f(p(c_{1}),...,p(c_{|\mathbb{A}|});\alpha_{1},...,\alpha_{|\mathbb{A}|})=\frac{1}{\mathbf{B}(\mathbf{a})}\prod_{i=1}^{|\mathbb{A}|}P(c_{i})^{\alpha_{i}-1}$, $\alpha_{i}=1\forall i$ and assigned to each object in the inventory. Similarly, a probability is sampled from a flat Dirichlet to assign the occurrence probability of the objects within each category. More details can be found in the Appendix 1.
@@ 3 Learning abstractions from sequences
A fundamental feature of the model is to parse sequences in chunks as the basic cognitive units (Miller, 1956). Along with each parse $t$, the transition counts between the previous $c_{L}$ and next chunks $c_{R}$ are recursively updated: $T_{ij}(t+1)=T_{ij}(t)+[i=c_{L}][j=c_{R}]$ ($[\cdot]=1$ if the argument is true and 0 otherwise) and along with the identification frequency of each parsed chunk $M_{i}(t+1)=M_{i}(t)+[i=c]$. When modeling human behavior, each entry of $M$ and $T$ multiplies with a memory decay parameter $\theta$ per parsing step. The probability of observing a sequence of parsed chunks $c_{1},c_{2},...,c_{N}$ the becomes $P(c_{1},c_{2},...,c_{N})=\prod_{c_{i}\in\mathbb{C}}P_{\mathbb{C}}(c_{i})$.
@@
Amongst the correlated consecutive chunk pairs, the model identifies all eligible transitions by intersecting the post-adjacency columns with the pre-adjacency rows, proposing a set of chunks $E(v)={c_{1},c_{2},...,c_{j}}$ represented by a new variable $v$. This variable proposal is accepted if the set of chunks that satisfy the condition $T_{min}\leq|E(v)|\leq T_{max}$ and $\sum_{c_{i}}M(c_{i})\geq freq_{T}$. A variable denotes a set of chunks $E(v)$ and is identified if any of the chunks that it denotes is parsed (Figure 2c). Together, the common preceding chunk, in conjunction with the newly proposed variable, followed by the common succeeding chunk, is concatenated and proposed as a novel chunk $c_{L}\oplus v\oplus c_{R}$ to add to the inventory $\mathbb{C}$. An edge that connects $c_{L}$ to the new chunk $c_{L}\oplus v\oplus c_{R}$ to be added to the parsing graph, as illustrated in Figure 2d. This process introduces chunks with embedded variables and takes the computational complexity of $O(PQ)$, where $P$ and $Q$ are the numbers of unique entries of the adjacency and postadjaency chunks amongst the correlated consecutive chunk pairs. Two variables are merged into one if they share the same preceding and succeeding chunks. During parsing, a chunk is consistent with the sequence if any of its included variables contain a denoting chunk that is consistent with the sequence. Upon identifying a chunk during parsing, the marginal and transition count for both the chunk and the immediate variables that the chunk belongs to is incremented.
@@ 4.2 Learning from real-world sequential data
Going beyond artificially generated sequences, we evaluate HVM, HCM, and LZ78 in text domains from the BabyLM language dataset, which contain text snippets from a collection of data domains Warstadt et al. (2023). For each data domain, we took random snippets of 1000 characters and calculated evaluation metrics, including compression ratio (the number of tokens before compression divided by the number of tokens after compression), sequence complexity (negative log-likelihood), and compression efficiency (the length of the compressed sequence divided by the number of dictionary entries).
@@ 4.4 Comparing abstraction learning and transfer in LLMs
Given that large language models (LLMs) have been praised for their emergent abilities (Wei et al., 2022), yet also compared to "lossy compression" (Chiang, 2023) and struggle with tasks that demand abstract thought (Fleuret et al., 2011; Odouard & Mitchell, 2022), we want to investigate whether models with or without abstraction better describe them. Specifically, we ask whether LLMs abstract similarly to humans and how their behavior compares to cognitive models incorporating abstraction. To explore this question, we conduct the same sequence recall experiment on LLMs.
To do that, we adapt the instructions from the human experiment into prompts for in-context learning in LLMs. The LLMs are tasked with predicting the next subsequent tokens based on the context of previously seen instruction sequences. The specific prompt is taken from the instruction given to the human participants: “Please remember the following sequence: BADFDCBFFEDB, …, BCDFDCBFFADB.” We then calculate the conditional probability of the next token, $z_{i}$, given the instruction sequence history up until the current token $P(z_{i}|prompt)$. After obtaining this probability, the subsequent token in the instruction is added to the prompt, and the LLM’s conditional probability for the following token is updated accordingly. Using this prompt-chaining method, we can calculate LLM’s negative log-likelihood for the next tokenized instruction sequence: $-\log P(S)=-\log P(z_{1}|prompt)\prod_{i=2,...,n}P(z_{i}|prompt_{i-1})$, analogous to the cognitive models. We apply this approach to evaluate the token prediction likelihood in four large language models: GPT-2 (Radford et al., 2019) and three variations of the Llama 2 model (Touvron et al., 2023)—Llama with 7 billion parameters (Llama 7B), 13 billion parameters (Llama 13B), and 70 billion parameters (Llama 70B).
@@
4.5 Abstraction, distortion, generalization
@@
Figure 6: a. New chunks or variables are created from previously learned ones, adding increasingly nested structures to the representation graph. b. As the allowable distortion increases (moving right on the x-axis), the required rate to encode the data decreases (moving down on the y-axis). Introducing variables enables a more compressed sequence, which trades off with higher uncertainty introduced by the variables. c. Abstraction and transfer: The higher the abstraction layer, the higher the likelihood (lower negative log-likelihood) that HVM parses novel sequences.
With a model that can learn interpretable abstraction, we can delve into the relationship between learning layers of abstraction, compression, and uncertainty. HVM updates its dictionary each iteration by proposing chunks and variables, which are used to parse the sequence for the next iteration. New chunks or variables are created from previously learned ones, adding increasingly nested structures to the representation graph. This feature affords a controlled accessment of the relation between the level of abstraction, the amount of distortion that abstract representation introduces, and its relation to transfer. We measure nestedness using the representation complexity $RC(\mathbb{V})=\sum_{v\in\mathbb{V}}\sum_{u\in\mathbb{E}(v)}-\log P(u|v)$, which is closely related to the encoding cost of the representation graph A.5. We also measured the uncertainty carried by the embedded variables in the chunks via entropy formulated as $\sum_{c\in\mathbb{C}}P(c)\sum_{v\in V(c)}\sum_{u\in E(v)}-P(u|v)\log P(u|v)$.
@@
Figure 6a shows the relation between the three factors as a consequence of increasing abstraction learning iterations. As the representation graph becomes increasingly nested, the newly learned chunks explain a longer part of the sequence, and the chunks to parse the sequence contain more uncertain variables. This trade-off between compression and uncertainty reflects rate-distortion theory specifying that the best possible compression of a signal $X$ contains a lower bound on quality loss specified by the Rate-Distortion Function (R(D)): $R(D)=\inf_{Q}R_{Q}\text{s.t.}\mathbb{E}[d(X,\hat{X})]\leq D$ (Shannon, 1959; Cover & Thomas, 2012). Taking representation complexity as a distortion function, as the level of abstraction increases, HVM learns patterns that span and predict longer parts of the sequence while more distortion and uncertainty are introduced. This observation on distortion with the level of abstraction resonates with previous findings on an increased level of mental errors with learning more abstract representations in humans (Lynn et al., 2020).
@@
Higher levels of abstraction implies more flexible transfer If learning higher abstraction layers implies more distortion, what would be the benefit apart from compression? We suggest that the other side of the coin lies in generalization and demonstrates the effect of abstraction on model transfer. We trained HVMs with increasing abstraction layers and evaluated the models’ transfer likelihood upon parsing a novel sequence. As shown in Figure 6b, the higher the abstraction layer, the higher the likelihood (lower negative log-likelihood) that HVM parses the transfer sequence. To compare, we also evaluated the transfer performance of the chunking model (HCM) with an increasing level of iteration. While the transfer performance for HCM improves with more layers of chunk learning, it converges at a higher negative log-likelihood. Learning an increasing abstract representation by the HVM enables it to compress novel transfer sequences in a more succinct way. Furthermore, the more abstraction layers are acquired, the higher the relative advantage for HVM that learns abstraction holds over the HCM that only learns chunks.

### v2

@@ Abstract
Humans excel at learning abstract patterns across different sequences, filtering out irrelevant details, and transferring these generalized concepts to new sequences. In contrast, many sequence learning models lack the ability to abstract, which leads to memory inefficiency and poor transfer. We introduce a non-parametric hierarchical variable learning model (HVM) that learns chunks from sequences and abstracts contextually similar chunks as variables. HVM efficiently organizes memory while uncovering abstractions, leading to compact sequence representations. When learning on language datasets such as babyLM, HVM learns a more efficient dictionary than standard compression algorithms such as Lempel-Ziv. In a sequence recall task requiring the acquisition and transfer of variables embedded in sequences, we demonstrate HVM’s sequence likelihood correlates with human recall times. In contrast, large language models (LLMs) struggle to transfer abstract variables as effectively as humans. From HVM’s adjustable layer of abstraction, we demonstrate that the model realizes a precise trade-off between compression and generalization. Our work offers a cognitive model that captures the learning and transfer of abstract representations in human cognition and differentiates itself from LLMs.
@@ 1 Introduction
We hypothesize that the world contains patterns across scales of time and abstraction. Intelligent agents-facing sequences with nested hierarchical structures- need to model these structures to store, process, and interact in such environments. As a rational strategy, intelligent agents shall characterize the temporal structure via chunking and characterize the abstract structure via identifying different items that play similar roles. To explore these operations, we design a generative model that produces sequences nested with hierarchical structure and propose an approximate recognition model that conjunctively learns to chunk and abstract.
@@
We go beyond previous proposals that introduce chunking as a mechanism for learning to compose complex structures from elementary perceptual units (Gobet et al., 2001; Miller, 1956; Wu et al., 2022) and propose a cognitive model that combines chunking and abstraction in one single system. The model uses abstraction in two key ways: first, by identifying shared features to facilitate efficient pattern retrieval; and second, by categorizing different sequential items that appear in the same context, much like a variable in a computer program. These paired mechanisms enable the model to parse sequences into chunks and form abstract patterns based on both concrete and previously learned abstractions, allowing for compact representations while revealing recurring patterns at the level of abstract categories. Consequentially, increasingly complex abstract patterns can be discovered layer by layer.
@@
We first demonstrate the benefits of abstraction in memory efficiency and sequence parsing by comparing our algorithm with previous chunking models and other dictionary-based compression methods. Then, we show that the model exhibits human-like signatures of abstraction in a memory experiment requiring the transfer of abstract concepts. In the same experiment, we contrast the model’s generalization behavior with large language models (LLMs). Additionally, we demonstrate the connection between abstraction level and abstract concept transfer by varying the level of abstraction as a parameter in the model. Our work offers a cognitive model that captures the learning and transfer of abstract representations in human cognition and differentiates itself from the behavior of artificial agents.
@@
Figure 1: Generative Model. The inventory is initialized with a set of atomic units, which randomly combine into objects. Some objects are randomly selected to become categories. Categories of existing objects have similar interaction properties. The dashed arrow points from those objects to their corresponding categories. These categories further combine with existing objects. An observational sequence is composed of objects randomly sampled from the existing inventory. Categories are latent and never explicitly observed but are manifested in one of the denoting objects.
@@ 2 Generating sequences with objects that contain hierarchical abstract structures
Hierarchical structures have been long presumed to characterize natural sequences — such as molecules composed of chemical elements or the hierarchical nature of language (Abler, 1989; Sportiche et al., 2013; von Humboldt & Losonsky, 1999). To study models respect hierarchical structure, we take a standard approach starting from a conformative generative model (Teh, 2006; Milan et al., ). We design a probabilistic generative model that generates sequences by sampling objects from an inventory which contains a hierarchical structure.
The generative model creates an inventory of recurring objects over $d$ iterations of expansion. As illustrated in Figure 1, the inventory starts with a set of atomic units $\mathbb{A}$. On each iteration, a novel object or category is created equiprobably. A new category (graded node) is created by pointing to a random selection of objects from the inventory up to the moment (objects are treated disjunctively). A new object is created by concatenating a random selection of pre-existing objects or categories from the existing inventory. After the inventory has been expanded up to the $d$-th iteration, the independent occurrence probability in the sequence is sampled from a flat Dirichlet distribution $f(p(c_{1}),...,p(c_{|\mathbb{A}|});\alpha_{1},...,\alpha_{|\mathbb{A}|})=\frac{1}{\mathbf{B}(\mathbf{a})}\prod_{i=1}^{|\mathbb{A}|}P(c_{i})^{\alpha_{i}-1}$, $\alpha_{i}=1\forall i$ and assigned to each object in the inventory. Similarly, a probability is sampled from a flat Dirichlet to assign the occurrence probability of the objects within each category. More details can be found in the Appendix 1.
@@ 3 Learning abstractions from sequences
A fundamental feature of the model is to parse sequences in chunks as the basic cognitive units (Miller, 1956). Along with each parse $t$, the transition counts between the previous $c_{L}$ and next chunks $c_{R}$ are recursively updated: $T_{ij}(t+1)=T_{ij}(t)+[i=c_{L}][j=c_{R}]$ ($[\cdot]=1$ if the argument is true and 0 otherwise) and along with the identification frequency of each parsed chunk $M_{i}(t+1)=M_{i}(t)+[i=c]$. When modeling human behavior, each entry of $M$ and $T$ multiplies with a memory decay parameter $\theta$ per parsing step. The probability of observing a sequence of parsed chunks $c_{1},c_{2},...,c_{N}$ then becomes $P(c_{1},c_{2},...,c_{N})=\prod_{c_{i}\in\mathbb{C}}P_{\mathbb{C}}(c_{i})$.
@@
Among the correlated consecutive chunk pairs, the model identifies all eligible transitions by intersecting the post-adjacency columns with the pre-adjacency rows, proposing a set of chunks $E(v)={c_{1},c_{2},...,c_{j}}$ represented by a new variable $v$. This variable proposal is accepted if the set of chunks that satisfy the condition $T_{min}\leq|E(v)|\leq T_{max}$ and $\sum_{c_{i}}M(c_{i})\geq freq_{T}$. A variable denotes a set of chunks $E(v)$ and is identified if any of the chunks that it denotes is parsed (Figure 2c). Together, the common preceding chunk, in conjunction with the newly proposed variable, followed by the common succeeding chunk, is concatenated and proposed as a novel chunk $c_{L}\oplus v\oplus c_{R}$ to add to the inventory $\mathbb{C}$. An edge that connects $c_{L}$ to the new chunk $c_{L}\oplus v\oplus c_{R}$ to be added to the parsing graph, as illustrated in Figure 2d. Two variables are merged into one if they share the same preceding and succeeding chunks. During parsing, a chunk is consistent with the sequence if any of its included variables contain a denoting chunk that is consistent with the sequence. Upon identifying a chunk during parsing, the marginal and transition count for both the chunk and the immediate variables that the chunk belongs to is incremented.
@@ 4.2 Learning from real-world sequential data
Going beyond artificially generated sequences, we evaluate HVM, HCM, and LZ78 in text domains from the BabyLM language dataset, which contain text snippets from a collection of data domains such as conversational child-directed speech, narrative stories, and other linguistically rich environments Warstadt et al. (2023). For each data domain, we took random snippets of 1000 characters and calculated evaluation metrics, including compression ratio (the number of tokens before compression divided by the number of tokens after compression), sequence complexity (negative log-likelihood), and compression efficiency (the length of the compressed sequence divided by the number of dictionary entries).
@@ 4.4 Comparing abstraction learning and transfer in LLMs
In recent years, large language models (LLMs) have demonstrated their emergent abilities (Wei et al., 2022), raising the concern that their intelligence will soon exceed that of humans, while at the same time, LLMs exert severe limitations in a set of tasks that demands abstract thought (Fleuret et al., 2011; Odouard & Mitchell, 2022). Given the critical role that abstraction plays in reasoning and generalization, it becomes increasingly important to delineate the aspects of LLM that most significantly deviate from human cognition, and find out what features inside a cognitive model would provide a missing behavioral aspect inside LLMs. The missing feature will inform the limitations of LLM’s computational component that is predictive of its ability to generalize and transfer. Specifically, we ask whether LLMs abstract similarly to humans and how their behavior compares to cognitive models incorporating abstraction. To explore this question, we conduct the same sequence recall experiment on LLMs.
We adapt the instructions from the human experiment into prompts for in-context learning in LLMs. The LLMs are tasked with predicting the next subsequent tokens based on the context of previously seen instruction sequences. The specific prompt is taken from the instruction given to the human participants: “Please remember the following sequence: BADFDCBFFEDB, …, BCDFDCBFFADB.” We then calculate the conditional probability of the next token, $z_{i}$, given the instruction sequence history up until the current token $P(z_{i}|prompt)$. After obtaining this probability, the subsequent token in the instruction is added to the prompt, and the LLM’s conditional probability for the following token is updated accordingly. Using this prompt-chaining method, we can calculate LLM’s negative log-likelihood for the next tokenized instruction sequence: $-\log P(S)=-\log P(z_{1}|prompt)\prod_{i=2,...,n}P(z_{i}|prompt_{i-1})$, analogous to the cognitive models. We apply this approach to evaluate the token prediction likelihood in four large language models: GPT-2 (Radford et al., 2019) and three variations of the Llama 2 model (Touvron et al., 2023) and two variations of the Llama 3 model.
@@
4.5 Abstraction, distortion, and generalization
@@
Figure 6: a. New chunks or variables are created from previously learned ones, adding increasingly nested structures to the representation graph. b. As the allowable distortion increases (moving right on the x-axis), the required rate to encode the data in terms of chunks decreases (moving down on the y-axis). Introducing variables enables a more compressed sequence, which trades off with higher representational complexity introduced by the variables embedded inside the chunks. c. Abstraction and transfer: The higher the abstraction layer, the higher the likelihood (lower negative log-likelihood) that HVM parses novel sequences.
@@
With a model that can learn interpretable abstraction, we can delve into the relationship between learning layers of abstraction, compression, and uncertainty. HVM updates its dictionary each iteration by proposing chunks and variables, which are used to parse the sequence for the next iteration. New chunks or variables are created from previously learned ones, adding increasingly nested structures to the representation graph. This feature affords a controlled assessment of the relation between the level of abstraction, the amount of distortion that abstract representation introduces, and its relation to transfer and generalization to novel sequences. We measure nestedness using the representation complexity $RC(\mathbb{V})=\sum_{v\in\mathbb{V}}\sum_{u\in\mathbb{E}(v)}-\log P(u|v)$, which is closely related to the encoding cost of the representation graph A.5. We also measured the uncertainty carried by the embedded variables in the chunks via entropy formulated as $\sum_{c\in\mathbb{C}}P(c)\sum_{v\in V(c)}\sum_{u\in E(v)}-P(u|v)\log P(u|v)$. Figure 6a shows the relation between the three factors as a consequence of increasing abstraction learning iterations. As the representation graph becomes increasingly nested, the newly learned chunks explain a longer part of the sequence, and these chunks embed more uncertain variables. This trade-off between compression and uncertainty reflects rate-distortion theory specifying that the best possible compression of a signal $X$ contains a lower bound on quality loss specified by the Rate-Distortion Function (R(D)): $R(D)=\inf_{Q}R_{Q}\text{s.t.}\mathbb{E}[d(X,\hat{X})]\leq D$ (Shannon, 1959; Cover & Thomas, 2012). Taking representation complexity as a distortion function, as the level of abstraction increases, HVM learns patterns that span and predict longer parts of the sequence while more distortion and uncertainty are introduced. This observation on distortion with the level of abstraction resonates with previous findings relating mental errors with learning more abstract representations in humans (Lynn et al., 2020).
@@
Higher levels of abstraction implies more flexible transfer If learning higher abstraction layers implies more distortion, what would be the benefit apart from compression? We suggest that the other side of the coin lies in generalization and demonstrates the effect of abstraction on model transfer. We trained HVMs with increasing abstraction layers and evaluated the models’ transfer likelihood upon parsing a novel sequence. As shown in Figure 6b, the higher the abstraction layer, the higher the likelihood (lower negative log-likelihood) that HVM parses the transfer sequence. To compare, we also evaluated the transfer performance of the chunking model (HCM) with an increasing level of abstraction iteration. While the transfer performance for HCM improves with more layers of chunk learning, it converges at a higher negative log-likelihood. Learning an increasing abstract representation by the HVM enables it to compress novel transfer sequences in a more succinct way. Furthermore, the more abstraction layers are acquired, the higher the relative advantage for HVM that learns abstraction holds over the HCM that does not learn variables.